Consider a consumer of goods A, B, and C. PA =2, PB=6 and PC=10, on the basis of the hypothetical data given below, determine the optimum of the consumer when the income of the consumer is birr 130.


Quantity demand

MUA

MUB

MUC


1

24

120

140


2

22

110

120


3

20

96

100


4

18

80

100


5

16

64

60


6

14

48

50


7

12

42

36


8

10

36

20


9

8

30

6


10

6

24

2


Based on the above given find: -

A.The optimum combination of quantity

B.Calculate total utility (TU) at optimum combinations





1
Expert's answer
2021-05-04T12:16:23-0400

Solution:

A.). The optimum combination of quantity:



The optimum combination of quantity are as follows:

Good A = 8

Good B = 9

Good C = 6

This can be verified as follows:

Total income = (PA x 8) + (PB x 9) + (PC x 6)

130 = (2 "\times" 8) + (6 "\times" 9) + (10 "\times" 6)

130 = 16 + 54 + 60

130 = 130

 

B.). Calculate Total Utility (TU) at optimum combinations:

TU = 136 + 626 + 570 = 1,332

Total Utility (TU) at optimum combinations = 1,332